### v1

# Interactive, and Editable Simulation for Autonomous Driving

Abstract
Simulation frameworks have been key enablers for the development and validation of autonomous driving systems. However, existing methods struggle to comprehensively address the autonomy-oriented requirements of balancing: (i) dynamical fidelity, (ii) photorealistic rendering, (iii) context-relevant scenario orchestration, and (iv) real-time performance. To address these limitations, we present a unified framework for creating and curating high-fidelity digital twins to accelerate advancements in autonomous driving research. Our framework leverages a mix of physics-based and data-driven techniques for developing and simulating digital twins of autonomous vehicles and their operating environments. It is capable of reconstructing real-world scenes and assets (real2sim) with geometric and photorealistic accuracy and infusing them with various physical properties to enable real-time dynamical simulation of the ensuing driving scenarios. Additionally, it also incorporates a large language model (LLM) interface to flexibly edit the driving scenarios online via natural language prompts. We analyze the presented framework in terms of its fidelity, performance, and serviceability. Results indicate that our framework can reconstruct 3D scenes and assets with up to 97% structural similarity, while maintaining frame rates above 60 Hz. We also demonstrate that it can handle natural language prompts to generate diverse driving scenarios with up to 95% repeatability and 85% generalizability.
Index Terms:
Agentic AI, Generative AI, Spatial AI, Scene Reconstruction, Scenario Generation, Autonomous Vehicles, Digital Twins, Real2Sim Transfer
I Introduction
Figure 1: Proposed framework for autonomy-oriented digital twins, which blends photorealism and physical interaction with LLM-driven scenario orchestration for enhanced serviceability. Video: https://youtu.be/ZYlcV2J1MME
The development and deployment of autonomous driving systems requires extensive training/optimization and testing/validation in a wide variety of scenarios to ensure safety, robustness, and scalability. Real-world training/testing is often limited by high costs, large time investments, safety concerns, and the lack of control over creating edge cases or extreme conditions. In this context, digital twins offer a tantalizing alternative by providing controllable, repeatable, and diverse simulations for accelerating synthetic data generation, design optimization as well as systematic verification and validation. However, creating high-fidelity digital twins that simultaneously balance photorealism, physical accuracy, operational flexibility, and real-time performance remains a significant challenge.
To this end, we propose a unified framework (refer Fig. 1) for reconstructing high-fidelity digital twins and reconfiguring them via artificially intelligent (AI) agents. Our framework is designed to meet the specific requirements of autonomy-oriented digital twins: (a) physical accuracy and interactivity to simulate vehicle dynamics, sensor characteristics, and environment physics; (b) geometric and visual fidelity to support feature-rich and reliable perception, (c) contextual understanding for generating diverse driving scenarios, and (d) real-time interfacing with autonomous driving software stacks.
The core novelty of the framework lies in its ability to reconstruct real-world scenes and assets with geometric and photorealistic accuracy and seamlessly blend them with synthetic 3D elements represented by point, surface, or mesh models through a real-time hybrid rendering pipeline. By infusing physical properties into these assets, the framework enables real-time dynamical simulation of driving scenarios, capturing realistic interactions between vehicles, pedestrians, and environmental elements. To further enhance its serviceability, the framework integrates an intuitive human-machine interface (HMI) mediated by large language models (LLMs), which allows users to perform creative and interesting edits to generate a diverse range of driving scenarios in real-time via natural language prompts.
Unlike existing solutions that often focus on isolated aspects like physics (e.g., vehicle dynamics), graphics (e.g., static scene rendering), or context (e.g., predefined scenarios), our framework offers a comprehensive reality-to-simulation (real2sim) workflow with real-time performance. This holistic approach not only enhances simulation fidelity but also enables convenient prototyping of autonomous driving algorithms.
The key contributions of this paper are summarized below:
  • High-Fidelity Reconstruction: We provide a method to reconstruct autonomy-oriented scenes and assets from camera and LIDAR data and seamlessly create visually, geometrically, and physically accurate, interactive simulations. This allows data-driven digital twins to co-exist with physics-based digital twins in the same scene.
  • Automated Reconfiguration: We present a hierarchical method comprising a mix of LLM agent(s) as well as rule-based agent(s) to automatically generate driving scenarios based on natural language prompts. The proposed method fosters the creativity of generative AI, while enforcing rule-based bounds to ensure the generation of diverse yet pragmatic scenarios.
  • Experimental Evaluation: We experiment with the reconstruction and reconfiguration techniques to analyze their accuracy, performance, repeatability, and generalizability. We also present qualitative insights in terms of data expectations and prompt engineering.
  • Open-Source Contribution: We openly release the proposed framework as a part of AutoDRIVE Ecosystem [^1] AutoDRIVE: https://autodrive-ecosystem.github.io [1], to enable further research in the field of autonomous driving and digital twins.
The remainder of this paper is structured as follows: Section II discusses related work and positions our framework within the existing landscape. Section III details the architecture and implementation of the framework, including its core subsystems. Section IV presents experimental results and performance benchmarks. Finally, Section V concludes with key insights and future directions.
II Related Work
Simulation frameworks play a crucial role in the field of autonomous driving, as they alleviate the monetary, safety, spatial, and temporal constraints imposed during physical development and validation. However, state-of-the-art frameworks struggle to deliver a unified solution that combines truly photorealistic rendering with physically accurate, real-time simulations. This includes the system-of-systems level simulation of vehicles (sensors, actuators, vehicle dynamics), environments (structure, appearance, variability), and their complex interactions. Additionally, most of the existing approaches lack efficient user interfaces to orchestrate context-aware scenarios that can capture the diverse nature of real-world driving conditions. Below, we review some of the related works, which can be broadly categorized based on their isolated focus on scene reconstruction, scenario reconfiguration, or autonomy-oriented simulation.
II-A Scene Reconstruction
Figure 2: Proposed approach to (a) photorealistic and geometrically accurate 3D scene reconstruction, (b) hybrid scene composition for physically interactive and graphically realistic simulation, and (c) LLM-guided context-aware scenario reconfiguration.
Scene reconstruction methods vary in complexity, accuracy, and application. Each method has unique strengths, with trade-offs between quality, speed, and computational cost. Manual hand-crafting has been shown to achieve good results in some cases, but it is severely time-consuming, labor-intensive, and requires significant artistic skill. Additionally, manual reconstruction often leaves inaccuracies, which impact the environment representation and sensor simulation. Surface reconstruction methods [2] generate continuous and water-tight 3D surfaces from sparse point clouds, excelling in accuracy, but can struggle with noisy or incomplete data, and generally do not capture the visual appearance. Photogrammetry techniques [3, 4] efficiently synthesize 3D models via images captured from multiple viewpoints by analyzing the overlap and spatial relationships between the images, delivering close geometric and visual resemblance, though they require precise image capture and may struggle with the reconstruction of large-scale assets and scenes. Neural radiance fields (NeRFs) [5, 6] leverage modern deep learning methods to implicitly learn an underlying continuous volumetric scene function using a sparse set of input views to generate 3D photorealistic reconstructions by synthesizing views from different perspectives. This method excels at reconstructing complex geometries and textures but demands high computational resources and is significantly slower in terms of rendering. 3D Gaussian splatting (3DGS) [7, 8] offers a fast and memory-efficient solution for scene reconstruction by employing light-weight 3D Gaussian functions to represent points in space as splats parametrized by their pose, scale, color, and opacity, which collectively contribute to the appearance and geometry of reconstruction. This method often outperforms traditional voxel-based methods in terms of speed and quality, specifically excelling in the reconstruction of large-scale scenes and assets, which are central to autonomous driving simulations. However, there is no “one size fits all” solution to scene/asset reconstruction. Hence, we adopt a mix of some of the above techniques such that they complement each other’s strengths and weaknesses.
II-B Scenario Reconfiguration
Scenario reconfiguration methods offer different approaches to generating or modifying scenarios, each with its strengths and limitations, balancing between customization, automation, flexibility, and computational efficiency. Manual design provides full control over the design process, allowing for highly customized and specific scenarios, but it is time-consuming, labor-intensive, and requires expert skill. Data-driven methods [9] leverage large datasets to automatically generate or modify scenarios, offering scalability and efficiency, but are limited by the quality and diversity of the data used, potentially producing generic or unrealistic results. Adversarial generation techniques [10, 11] use generative models and search algorithms to create or transform scenarios into safety-critical corner or edge cases, but often require significant computational power and are sensitive to sampled data quality. Knowledge-based generation [12] utilizes rules, heuristics, or domain-specific information to generate scenarios, ensuring consistency and relevance, but can be less flexible and may struggle with unexpected or novel scenarios. Overall, most of the aforementioned methods have quite steep learning curves and are very systematic, and therefore limited in terms of creative exploration of novel scenarios. Thus, in our work, we propose a hierarchical method comprising a mix of LLM agent(s) and rule-based agent(s) to generate driving scenarios on demand through user-convenient natural language prompts. The proposed method fosters the creativity of generative AI (LLMs can come up with really interesting suggestions, sometimes even beyond the user’s imagination) while enforcing rule-based bounds to ensure the generation of diverse yet pragmatic scenarios.
II-C Autonomy-Oriented Simulation
Simulation frameworks vary in terms of their focus and fidelity. In the context of autonomous driving, the foci can be broadly categorized into dynamics and graphics. On one end of this spectrum, certain frameworks (e.g., [13, 14]) stress vehicle dynamics or terramechanics simulations from multi-body models to finite element analyses. Such frameworks usually compromise on visual fidelity or real-time performance. On the other end of the spectrum, some frameworks (e.g., [15, 16]) only focus on photorealistic rendering, and claim to achieve high fidelity through real-world data replays or graphical animations within reconstructed scenes. Such methods may not be real-time and are sometimes run offline. Finally, there are some frameworks (e.g., [17, 18]) between these two extremes that try to strike a balance between dynamics and graphics. These methods generally use simplified dynamics representations coupled with physically-based rendering (PBR) to compose driving scenarios with hand-crafted digital assets. Another challenge lies in the fact that such simulated scenarios (static scenes or dynamic assets) are often not representative of any real-world entities, which degrades the simulation fidelity. All in all, none of the existing simulation frameworks supports truly photorealistic rendering coupled with hierarchically consistent dynamics and physical accuracy, while offering extremely intuitive and user-friendly scenario reconfiguration and orchestration. Our research aims to address this issue by proposing an openly accessible, high-fidelity yet real-time framework for democratizing AI-driven digital twins in autonomous driving by fusing literal photorealism and dynamical simulation with LLM-guided scenario editing.
III Research Methodology
III-A Scene Reconstruction
Figure 3: High-fidelity reconstruction: Digital twin of the CU-ICAR campus showing (a) photorealistic rendering via 3D Gaussian splatting; (b) geometric accuracy w.r.t. 3D LIDAR point cloud data; (c) dynamical realism via 3D surface reconstruction; (d) visualization of reconstructed RGB and depth channels in split halves; (e) co-existence of dynamic 3DGS assets like a passenger car, a traffic cone, and a pedestrian sign with 3DMM assets like a road barrier, cement rubble, and a pedestrian; (f) real-time autonomy-oriented simulation with sensor visualization.
The proposed scene/asset reconstruction pipeline (refer Fig. 2(a-b)) can leverage camera and LIDAR data from hand-held or vehicle-mounted sensors. Here, camera data primarily contributes to photorealism, and LIDAR data ensures geometric precision; both data contribute to hybrid scene composition.
The camera data is first preprocessed to extract frames from continuous video sequences, while ensuring consistent feature matching. We observed that camera resolution, exposure changes, and video frame rate have a significant impact on feature matching, as they directly affect image sharpness, brightness/contrast, and motion blur. The extracted frames are parsed through a structure-from-motion (SfM) pipeline to obtain camera intrinsics, undistorted images, reconstructed poses, and feature points. The feature points are used to initialize 3D Gaussians. The camera intrinsics and reconstructed poses are used to project the 3D Gaussians from those viewpoints and rasterize 2D images. These rasterized images are compared with the corresponding undistorted images to compute the photometric loss. The gradients flow back via a differentiable tile rasterizer to update the 3D Gaussians (pose, scale, color, opacity) by minimizing error between the rasterized output and ground truth images.
Particularly, 3D Gaussian Splatting (3DGS) [7] represents scenes and assets using a finite set of “splats”, denoted as ${\mathcal{G}}=\{\bm{g}\}$, where the intensity distribution of every splat follows a 3-dimensional Gaussian function. Each individual Gaussian, $\bm{g}({o},{\bm{\mu}},{\mathbf{q}},{\bm{s}},{\bm{c}})$ is parametrized by opacity ${o}\in(0,1)$, position ${\bm{\mu}}\in\mathbb{R}^{3}$ represented as Cartesian coordinates, orientation ${\mathbf{q}}\in\mathbb{R}^{4}$ represented as a quaternion, anisotropic scaling factors ${\bm{s}}\in\mathbb{R}^{3}_{+}$, and view-dependent color representation ${\bm{c}}\in\mathbb{R}^{F}$ expressed as coefficients of spherical harmonics. To compute the color $C$ of a rasterized pixel, $\mathcal{N}$ Gaussians that intersect with the pixel are sorted based on their distance to the camera center (sorted by index $i\in\mathcal{N}$) and blended using alpha compositing as:
$C=\sum_{i\in\mathcal{N}}{{\bm{c}}}_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (1)
where $\alpha_{i}={o}_{i}\cdot\exp\left(-\frac{1}{2}(\mathbf{p}-\bm{\mu}_{i})^{T}\bm{\Sigma}_{i}^{-1}(\mathbf{p}-\bm{\mu}_{i})\right)$ for point $\mathbf{p}$ with $\bm{\Sigma}_{i}$ representing the 2D projection covariance. The spatial information is captured by applying an affine transformation ${\mathbf{T}}=({\mathbf{R}},{\mathbf{t}})\in\mathbb{SE}(3)$ to all Gaussians in the set as:
${\mathbf{T}}\otimes{\mathcal{G}}={({o},{\mathbf{R}}{\bm{\mu}}+{\mathbf{t}},\mathrm{Rot}({\mathbf{R}},{\mathbf{q}}),{\bm{s}},{\bm{c}})}$ (2)
where $\mathrm{Rot}(\cdot)$ represents the rotation of a quaternion using rotation matrix.
The LIDAR data is first registered to generate a cohesive point cloud, which is then segmented to filter out any unwanted noise or stray artifacts from the data. Next, the point normals are estimated and aligned using a triangulation surface model. This post-processed (aligned) point cloud data (PCD) is used to reconstruct a 3D mesh model (3DMM) of the scene/asset.
Particularly, Poisson Surface Reconstruction (PSR) [2] recovers a 3D surface from oriented point clouds by solving a spatial Poisson equation. It treats aligned normals as samples of a vector field $\overrightarrow{V}$, seeking an implicit indicator function $\widetilde{\chi}$ whose gradient best matches $\overrightarrow{V}$ (i.e., $\nabla\widetilde{\chi}\approx\overrightarrow{V}$). By taking the divergence of the vector field formed by the normals, the problem reduces to solving the Poisson equation:
$\Delta\widetilde{\chi}\approx\nabla\overrightarrow{V}$ (3)
The solution yields an approximation of the indicator function $\widetilde{\chi}$ in a least-squares sense. The reconstructed surface is extracted as an isosurface of this indicator function. PSR ensures smooth, watertight surfaces and is particularly robust to noise and non-uniform sampling, making it well-suited for geometrically reconstructing real-world 3D scenes and assets.
The two reconstructed representations (3DGS and 3DMM) are fused to obtain “literally” photorealistic (refer Fig. 3(a)), geometrically accurate (refer Fig. 3(b) and 3(d)), and physically interactive (refer Fig. 3(c) and 3(e)) 3D scenes/assets. This fusion is enabled by a hybrid rendering pipeline comprising hierarchical passes for 3DGS rendering and physically-based rendering (PBR). Consequently, these reconstructed scenes and assets can co-exist with other (hand-crafted, procedurally generated, etc.) 3DGS or 3DMM assets and scenes (refer Fig. 3(e)), enriching scenario generation capabilities. As an example, Fig 3(f) demonstrates the reconstructed CU-ICAR campus (3DGS) within AutoDRIVE Simulator [19], simulating the autonomy-oriented digital twin of OpenCAV [^2] OpenCAV: https://sites.google.com/view/opencav [20] (3DMM) including vehicle dynamics, sensor physics, and environment interactions.
III-B Scenario Reconfiguration
The proposed scenario reconfiguration pipeline (refer Fig. 2(c)) can leverage a hierarchical architecture comprising LLM agent(s) as well as rule-based agent(s). The level-1 LLM agent (a.k.a. “Sr. Manager”) parses natural language inputs prompted by the human user(s) to determine the scenario design requirements. These requirements are then passed on to level-2 LLM agents (a.k.a. “Jr. Managers”) to determine the low-level tasks. These specialized tasks are then assigned to the appropriate rule-based agents (a.k.a. “Workers”) to modify the existing scene by searching, adding/removing/updating, or positioning/arranging the requested assets or incorporating specific properties/behaviors to them. This way, the proposed method fosters the creativity of generative AI, while enforcing rule-based bounds to ensure the generation of diverse yet pragmatic scenarios. It is to be noted that the agent hierarchy can be flexibly adjusted depending on the expected complexity and granularity of scenario reconfiguration.
Fig. 1 depicts the high-level aspects of the pipeline. The user inputs a natural language prompt within the designated text field of the graphical user interface (GUI). This prompt is then passed through a prompt engineering function, which adds relevant context and response format to the raw prompt. This ensures robust prompt handling and consistent response formatting for future stages of the hierarchy. As highlighted above, the prompt turns into requirements (processed by level-1 LLM agent), which then turn into specialized tasks (processed by level-2 LLM agents). These specific tasks are executed by rule-based agents to realize the requested scenario.
It is worth mentioning that scenario reconfiguration is an online real-time process, meaning that the simulator continues its functions of physics time-stepping, rendering, data transfer via application programming interfaces (APIs), accepting inputs and providing feedback via the graphical user interface (GUI) or the human-machine interface (HMI), etc. Additionally, being an online process, the pipeline supports multi-stage commands while preserving context from the previous prompts. This is especially useful when sequentially editing the scenario while accepting or rejecting previous changes.
IV Results and Discussion
IV-A Scene Reconstruction
The following metrics were used to assess the quality and performance of the scene/asset reconstructions:
PSNR: Peak signal-to-noise ratio (PSNR) quantifies the reconstruction fidelity by comparing the pixel-wise similarity between the rendered and the ground truth images. It focuses on pixel intensity differences and is derived from the mean squared error (MSE) or $L_{2}$ distance between the two images.
$\mathrm{PSNR}(x,y)=10\cdot log_{10}\left(\frac{\mathrm{MAX}_{I}^{2}}{\mathrm{MSE}(x,y)}\right)$ (4)
where, $\mathrm{MAX}_{I}$ is the maximum possible pixel value (e.g., $2^{\mathrm{n}}$ for $n$-bit images). A higher PSNR indicates better reconstruction quality (lower error). PSNR can measure even small pixel-level differences but may fail to capture perceptual distortions such as blurriness or loss of structural information.
TABLE I: Scene Reconstruction Benchmarking
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS $\uparrow$
$\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$
3DGS 23.71 2.07 0.78 0.09 0.20 0.09 63.55 23.84
NeRF 18.35 2.00 0.41 0.16 0.56 0.16 0.57 0.27
TABLE II: Scene Reconstruction Performance Evaluation
Scene/Asset Data (#) PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS $\uparrow$
CU-ICAR 1420 23.71 0.78 0.20 63.55
CGEC 1241 23.47 0.74 0.32 61.96
AuE Lab 545 35.10 0.97 0.17 69.36
Footpath (Lit) 70 32.20 0.93 0.09 66.09
Footpath (Unlit) 263 31.82 0.90 0.22 68.27
Passenger Car 971 24.93 0.88 0.19 79.19
Traffic Cone 530 35.11 0.95 0.06 62.65
Pedestrian Sign 1102 31.41 0.94 0.08 64.40
SSIM: Structural similarity index measure (SSIM) evaluates the perceptual structural similarity between the rendered image and the ground truth by comparing luminance $l(x,y)=\frac{2\mu_{x}\mu_{y}+c_{1}}{\mu_{x}^{2}+\mu_{y}^{2}+c_{1}}$, contrast $c(x,y)=\frac{2\sigma_{x}\sigma_{y}+c_{2}}{\sigma_{x}^{2}+\sigma_{y}^{2}+c_{2}}$, and structure $s(x,y)=\frac{\sigma_{xy}+0.5c_{2}}{\sigma_{x}\sigma_{y}+0.5c_{2}}$. It is designed to align with human visual perception.
$\mathrm{SSIM}(x,y)=[l(x,y)]^{\alpha}\cdot[c(x,y)]^{\beta}\cdot[s(x,y)]^{\gamma}$ (5)
where, $\{\alpha,\beta,\gamma\}>0$ weigh the relative importance of $l$, $c$, and $s$ respectively. SSIM ranges from -1 to 1, indicating the degree of correlation between the pair of images (1 means perfect similarity). It captures perceptual quality better than PSNR by focusing on local image structures, which are critical for evaluating how well a method reconstructs edges, textures, and scene consistency.
LPIPS: Learned perceptual image patch similarity (LPIPS) is a deep-learning-based perceptual metric that compares high-level features extracted from a neural network (e.g., AlexNet in our case) rather than pixel-wise differences. It measures how perceptually similar two images are by computing the $L_{2}$ distance between their feature activations at multiple layers of a pre-trained neural network.
$\mathrm{LPIPS}(x,y)=\sum_{l}\frac{1}{H_{l}\cdot W_{l}}\sum_{h,w}\\ \left\|w_{l}\odot\left(\phi^{l}(x)_{h,w}-\phi^{l}(y)_{h,w}\right)\right\|_{2}^{2}$ (6)
LPIPS score ranges from 0 (perfect similarity) to 1 (completely different). It evaluates perceptual quality at a high level, capturing differences in texture, details, and global image structure that are not detected by PSNR or SSIM.
FPS: Frames per second (FPS) is a performance measure to assess the rendering rate [^3] All experiments were conducted using a single laptop PC with 12th Gen Intel Core i9-12900H 2.50 GHz CPU, NVIDIA GeForce RTX 3080 Ti GPU, and 32.0 GB RAM.. It measures the number of frames $\mathrm{n_{frames}}$ rendered within a predefined time interval $T$.
$\mathrm{FPS}=\frac{\mathrm{n_{frames}}}{T}$ (7)
FPS values can range from 0 (no update) to upwards of 30 (real-time), with $>$60 FPS being a good benchmark for autonomy-oriented simulations.
We first benchmark two different photorealistic reconstruction methods, viz. NeRF and 3DGS. Table I presents the reconstruction results for the CU-ICAR scene (used as a benchmark). 3DGS outperforms NeRF by a large margin: not only are the quality metrics for 3DGS better than NeRF (PSNR +29.21%, SSIM +90.24%, LPIPS -64.29%), but 3DGS is over 100 times faster than NeRF for the same scene. The choice of 3DGS is thus justified, and further results will assume 3DGS reconstruction, unless mentioned otherwise.
Following the benchmarking experiment, we analyzed the performance of 3D reconstruction across various scenes and assets. Here, we considered scene variations in terms of their type (indoor/outdoor), scale (small/large), and lighting (lit/unlit). Diversity in assets was considered in terms of their materials (color, texture, opacity, albedo, etc.) and scale (small/large). Table II presents the data requirements, reconstruction quality, and rendering rate for 5 scenes and 3 assets.
It was observed that large outdoor scenes typically required upwards of 1k frames for effective reconstruction. Smaller (indoor) scenes, on the other hand, could be reconstructed with just half the data, but their depth estimation was observed to be slightly inaccurate, generating a few stray artifacts. This issue (inaccurate depth mapping leading to stray artifacts) was more prominently observed for the footpath scenes, where the data quantity was severely reduced. We also noticed that a well-lit footpath scene required significantly ($\sim$4$\times$) less data than a similar footpath scene, which was poorly lit.
Asset reconstructions also followed similar observations, where larger asset reconstructions were prone to reduced depth estimation accuracy, which could be attributed to data coverage limitations. For smaller assets, the data requirement was relatively less, but reflective (or retro-reflective) assets required slightly more data for improving depth estimation.
It was inferred that the scale or type of scene/asset did not directly govern the reconstruction fidelity. Instead, the quality (coverage, lighting, resolution, frame-rate, etc.) of the data mattered the most. However, it is also arguable that collecting good-quality data for smaller assets and indoor scenes is much easier than their larger and outdoor counterparts.
IV-B Scenario Reconfiguration
(a) Prompt: It’s daytime!
(b) Prompt: Make it look like night time.
(c) Prompt: Change weather to be foggy.
(d) Prompt: Make it rain.
(e) Prompt: Let it snow.
(f) Prompt: It’s a clear day.
(g) Prompt 1: Create cement rubble at the center of the scene.
(h) Prompt 2: Add traffic cones to mark the maintenance.
(i) Prompt 3: Also add road barriers around there.
(j) Prompt: Add a parked car at the JTEKT building entrance.
(k) Prompt: Add a mid-scale Ackermann-steered robot performing skidpad maneuver.
(l) Prompt: Make a jaywalking pedestrian cut across the ego vehicle’s path.
Figure 4: LLM-guided reconfiguration: Digital twin of the CU-ICAR campus being reconfigured to (a) daytime; (b) nighttime; (c) foggy weather; (d) rainfall; (e) snowfall; (f) clear weather; (g-i) sequentially add maintenance work assets; (j) include a parked car; (k) add a 1:5 scale mobile robot controlled to perform a specific maneuver; (l) include a jaywalking pedestrian.
The design of experiments to analyze the LLM-guided scenario reconfiguration pipeline was split into benchmarking (refer Table III) and evaluation (refer Table IV) categories.
In terms of benchmarking, we employed 3 different open-access LLMs, viz. Mistral 0.2, Llama 3.1, and Gemma 2. The choice of these models was primarily driven by their open accessibility, though it is worth mentioning that the proposed framework is capable of using any other models or APIs (both open as well as commercial). We consider the generalizability and repeatability of the model as our primary metrics for benchmarking, in addition to the model’s disk size occupancy, number of parameters, and inference time (from prompt input to scenario update). Here, repeatability is defined as the ability to produce the same result/response for 100 trials across 7 tasks with a direct prompt. Whereas, generalizability is defined as the ability to reconfigure the scenario satisfactorily for 100 trials across 7 tasks with 4 grades of prompts.
The prompt types ranged from explicit/direct to incorrect/erroneous, with a total of 4 different gradations as defined below:
  • Direct Prompt: These prompts directly express the user’s intent, with exact asset names, numeric values, and other data as applicable. Such commands can be expected from domain experts.
  • Indirect Prompt: These prompts clearly express the user’s intent but do not provide exact asset names, numeric values, or similar information. Such commands can be expected from users with a fair amount of knowledge, but who are not necessarily aware of specific terminology or conventions.
  • Vague Prompt: These prompts express the user’s intent minimally, and some things are left for interpretation by the model. Such commands can be expected from users with superficial or no knowledge.
  • Erroneous Prompt: These prompts try to express the user’s intent but include structural and/or grammatical errors. Such commands can be expected from non-native English speakers.
TABLE III: Scenario Reconfiguration Benchmarking
LLM Size (GB) Params. (#) Gen. (%) Rep. (%) Time (s)
Mistral 0.2 4.06 7B 46.21 92.65 6.60
Llama 3.1 4.58 8B 82.43 89.38 7.52
Gemma 2 5.36 9B 84.93 94.86 8.45
TABLE IV: Scenario Reconfiguration Performance Evaluation
Task/Satisfaction Direct Prompt Indirect Prompt Vague Prompt Erroneous Prompt
Search 99% 98% 80% 69%
Addition 99% 98% 90% 82%
Removal 100% 99% 88% 81%
Positioning 92% 86% 72% 65%
Moving 90% 87% 81% 73%
Arrangement 86% 85% 72% 57%
Appearance 98% 97% 79% 75%
The tasks included (i) searching for specific assets or properties, (ii) addition of assets, (iii) removal of assets, (iv) positioning the assets according to spatial cues, (v) moving the assets, (vi) arranging homogeneous or heterogeneous assets, and (vii) modifying appearance of the scene.
As expected, models with more parameters occupied higher disk space and incurred larger inference times. Particularly, it was observed that for every 1B additional parameters, the model size increased by $\sim$0.5 GB, and its inference time by $\sim$1 second. However, for the objective of scenario reconfiguration, each increment in the parameter space improved the model performance (generalizability and repeatability) significantly. This performance improvement outweighed the slight increase in model size, given modern digital storage systems. Similarly, a slightly higher inference time (order of milliseconds) could not be compared to a human user hand-crafting the scenario (order of hours or days).
In terms of model-specific performance, it was observed that Mistral 0.2 provided highly overconfident responses, resulting in high repeatability without much generalizability. Llama 3.1 and Gemma 2 performed comparably, but the latter was slightly better and generated more elegant responses. Consequently, further results will assume the Gemma 2 model, unless mentioned otherwise.
A detailed performance evaluation of the Gemma 2 model indicated that direct prompts resulted in over 90% satisfaction for the majority of the tasks, except for arrangement, which was an inherently complex task. Reducing the quality of prompts to indirect structuring resulted in degradation of positioning and moving tasks, without significantly affecting others. Further degrading the prompt quality to vague structuring resulted in a significant degradation of satisfaction across all the tasks, with search and appearance tasks being affected the most. Finally, adulterating the prompts with grammatical and/or structural errors resulted in a noticeable degradation of satisfaction across all the tasks, with arrangement and search tasks being affected the most.
We also conducted experiments with varying complexities of prompts and analyzed them qualitatively. These started with simple tasks like changing the time of day (refer Fig. 4(a-b)) or weather conditions (refer Fig. 4(c-f)). Next, we performed multi-stage scenario editing using sequential prompting (refer Fig. 4(g-i)) without repeating earlier information, to assess the context-retention ability. Finally, we also prompted complex commands involving several types of tasks to be executed simultaneously (refer Fig. 4(j-l)) to assess the task management ability. It was interesting to see the spatial awareness (e.g., implicit parking location in Fig. 4(j)) of the model, as well as its context-aware creativity (e.g., adding a pedestrian sign without being explicitly asked in Fig. 4(l)).
V Conclusion
This paper presented a novel, unified simulation framework designed to meet the multifaceted demands of autonomous driving research by combining dynamical fidelity, photorealistic rendering, scenario orchestration, and real-time performance. We demonstrated how the proposed framework effectively bridges the gap between real2sim fidelity and serviceability in digital twin simulations by integrating physics-based modeling with data-driven reconstruction and enabling natural language interaction through an LLM interface. Our experiments indicated high levels of structural similarity (97%) in reconstructed scenes/assets, strong generalizability (95%) and repeatability (85%) in scenario reconfiguration, and real-time performance ($>$60 FPS) on a consumer-grade PC. These findings suggest that the proposed framework can serve as a scalable tool for advancing the development, optimization, and validation of autonomous driving algorithms under realistic and diverse conditions.
Future research could focus on defining data requirements (quality and quantity) to further streamline the scene/asset reconstruction pipeline. Additionally, expanding the framework’s library of reconstructed scenes and assets would be a natural step towards enriching the creativity and variability of orchestrated simulations. Finally, there is a scope for enhancing the scenario reconfiguration pipeline by redefining the agentic AI topology, introducing additional context using techniques such as retrieval-augmented generation (RAG), or specializing it through fine-tuning.